MICROWAVE STUDY OF A HYDROGEN-TRANSFER-TRIGGERED METHYL-GROUP INTERNAL RO-TATION IN 5-METHYLTROPOLONE

VADIM V. ILYUSHIN, INSTITUTE OF RADIO ASTRONOMY OF NASU, CHERVONOPRA-PORNA 4, 61002 KHARKOV, UKRAINE; EMILY A. CLOESSNER, DEPARTMENT OF CHEM-ISTRY AND BIOCHEMISTRY, COLLEGE OF CHARLESTON, CHARLESTON, SC 29424, USA; YUNG-CHING CHOU, DEPARTMENT OF NATURAL SCIENCE, TAIPEI MUNICIPAL UNI-VERSITY OF EDUCATION, TAIPEI 10048, TAIWAN; LAURA B. PICRAUX, SUN CHEMICAL, CINCINNATI, OH 45232, USA; \underline{JON T. HOUGEN}, OPTICAL TECHNOLOGY DIVISION, NA-TIONAL INSTITUTE OF STANDARDS AND TECHNOLOGY, GAITHERSBURG, MD 20899-8441, USA; RICHARD LAVRICH, DEPARTMENT OF CHEMISTRY AND BIOCHEMISTRY, COLLEGE OF CHARLESTON, CHARLESTON, SC 29424, USA.

We present here the first experimental and theoretical study of the microwave spectrum of 5-methyltropolone, which can be visualized as a 7-membered aromatic carbon ring with a five-membered hydrogen-bonded cyclic structure at the top and a methyl group at the bottom. The molecule exhibits two large-amplitude motions, an intramolecular hydrogen transfer and a methyl torsion. The former motion is particularly interesting because transfer of the hydrogen atom from the hydroxyl to the carbonyl group induces a tautomerization in the molecule, which then triggers a 60^{∘} internal rotation of the methyl group. Measurements were carried out by Fourier-transform microwave spectroscopy in the 8 to 24 GHz frequency range. Theoretical analysis was carried out using a tunneling-rotational Hamiltonian based on a G_{12}^{m} extendedgroup-theory formalism. Our global fit of 1015 transitions to 20 molecular parameters gave a root-mean-square deviation of 1.5 kHz. The tunneling splitting of the two J = 0 levels arising from a hypothetical pure hydrogen transfer motion is calculated to be 1310 MHz. The tunneling splitting of the two J = 0 levels arising from a hypothetical pure methyl-top internal rotation motion is calculated to be 885 MHz. Some theoretical difficulties in interpreting the low-order tunneling parameters in this and the related molecule 2-methylmalonaldehyde will be discussed.